### v2

# Quasiregular maps of Sierpiński carpet Julia sets

Abstract.
We prove that if $f$ and $g$ are postcritically finite rational maps whose Julia sets $\mathcal{J}(f),\mathcal{J}(g)$, respectively, are Sierpiński carpets, and if $\xi$ is a quasiregular map of the Riemann sphere $\widehat{\mathbb{C}}$ with $\xi^{-1}(\mathcal{J}(g))=\mathcal{J}(f)$, then $\xi$ is the restriction of a rational map to the Julia set $\mathcal{J}(f)$. Moreover, when $g=f$ we prove that, for some positive integers $k$ and $l$, $f^{k}\circ\xi^{l}=f^{2k}$. These conclusions extend the main results of [BLM].
SM supported by NSF grant DMS-2247364.
Finally, we demonstrate that when Julia sets of postcritically finite rational maps are not Sierpiński carpets, say they are tree-like or gaskets, the above conclusions no longer hold.
1. Introduction
A Sierpiński carpet, or carpet for short, is a homeomorphic image of the standard Sierpiński carpet fractal. Due to [Wh], a subset $S$ of $\widehat{\mathbb{C}}$ is a Sierpiński carpet if and only if it has empty interior and can be written as
[…]
where $D_{k},\,k\in\mathbb{N}$, are pairwise disjoint Jordan domains with $\partial D_{k}\cap\partial D_{l}=\emptyset$ for $k\neq l$, and $\mathrm{diam}(D_{k})\to 0$ as $k\to\infty$. The boundaries $\partial D_{k}$ of $D_{k},\ k\in\mathbb{N}$, are called peripheral circles. A carpet $S$ is called a round Sierpiński carpet, or round carpet, if the domains $D_{k},\,k\in\mathbb{N}$, are open geometric disks.
A rational map $f$ is called postcritically finite if each critical point of $f$ has a finite forward orbit. The following result is Theorem 1.4 from [BLM].
Theorem 1.1 ([BLM]).
Let $f$ and $g$ be two postcritically finite rational maps and suppose that the corresponding Julia sets $\mathcal{J}(f)$ and $\mathcal{J}(g)$ are Sierpiński carpets. If $\xi$ is a quasisymmetric homeomorphism of $\mathcal{J}(f)$ onto $\mathcal{J}(g)$, then it is the restriction to $\mathcal{J}(f)$ of a Mobius transformation on $\widehat{\mathbb{C}}$.
Our first main result extends this theorem to quasiregular maps.
Theorem 1.2.
Let $f$ and $g$ be postcritically finite rational maps whose Julia sets $\mathcal{J}(f),\mathcal{J}(g)$, respectively, are Sierpiński carpets. If $\xi\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ is a quasiregular map such that $\xi^{-1}(\mathcal{J}(g))=\mathcal{J}(f)$, then $\xi|_{\mathcal{J}(f)}$ has a unique extension to a rational map $\widehat{\xi}\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ satisfying
(1) $G^{n+k}\circ\widehat{\xi}=G^{n}\circ\widehat{\xi}\circ F^{k},$
for all $k,n\in\mathbb{N},\ n\geq l$, for some $l\in\mathbb{N}$, where $F$ and $G$ are some iterates of $f$ and $g$, respectively.
Moreover, when $f=g$, we have the following result.
Theorem 1.3.
[…]
(2) $f^{k}\circ(\widehat{\xi})^{l}=f^{2k}.$
The structure of this paper is as follows. Section 2 introduces quasiregular and related maps. Section 3 is dedicated to the theory of Schottky maps, as developed by the first author [Me1, Me2, Me3], where the convergence Theorem 3.3 is an important ingredient of Proposition 5.1 in Section 5. Section 4 discusses the technique of conformal elevator from [BLM, Section 4], which enables us to construct local quasiconformal maps with images of definite size. This is another ingredient needed for Proposition 5.1. Section 5 outlines the proof of Proposition 5.1 adapted from [BLM] to a quasiregular case. After some preparation in Section 6 on rational maps sharing a measure of maximal entropy, based on [Le, Le2, LP, Ye], we prove the main Theorems 1.2 and 1.3 in Section 7. Section 8 discusses quasiregular maps between Julia sets of topologies other than the Sierpiński carpet. This section is inspired by the work of Y. Luo and D. Ntalampekos [LN], which reveals that gaskets do not exhibit the local rigidity properties of carpets. Our examples in this section add evidence to the contrast between carpet and gasket as well as tree-like Julia sets of polynomials.
Acknowledgements. The authors thank the Institute for Mathematical Sciences at Stony Brook University for its hospitality. The second author thanks the University of Science and Technology of China for the support of the travel that resulted in this work. The authors also thank Yusheng Luo for useful comments and references.
2. Quasiregular maps and related concepts
A non-constant mapping $f\colon U\to\widehat{\mathbb{C}}$ on a domain $U\subseteq\widehat{\mathbb{C}}$ is called $K$-quasiregular, $K\geq 1$, if $f$ is in the Sobolev space $W_{\mathrm{loc}}^{1,2}$ and for almost every $z\in U$,
$\left\lVert Df(z)\right\rVert^{2}\leq K\mathrm{det}(Df(z)).$
If $f$ is $K$-quasiregular for some $K\geq 1$, then $f$ is called quasiregular. An equivalent condition for $f\colon U\to\mathbb{C}$ to be quasiregular is that $f\in W_{\mathrm{loc}}^{1,2}$ and $f$ satisfies the Beltrami equation
$\frac{\partial f}{\partial\bar{z}}=\mu(z)\frac{\partial f}{\partial z}$
for some complex valued Lebesgue measurable function $\mu$ with
$\operatorname*{ess\,sup}|\mu|<1.$
If a $K$-quasiregular map $f$ is a homeomorphism onto its image in addition, then $f$ is called a $K$-quasiconformal map. A 1-quasiconformal map is called conformal.
We note here a useful factorization of quasiregular maps by S. Stoïlow [St] (see also [AIM, Section 5.5] and [LP2]):
Theorem 2.1 (Stoïlow’s Factorization).
If $f\colon U\to\widehat{\mathbb{C}}$ is a quasiregular map, then there exists a homeomorphism $h\colon U\to V$ which satisfies the same Beltrami equation as $f$. Moreover, for every such $h$ there exists a holomorphic map $g\colon V\to\widehat{\mathbb{C}}$ such that
$f=g\circ h.$
It follows easily from this theorem that a quasiregular map $f\colon U\to\widehat{\mathbb{C}}$ is a local homeomorphism at all but a discrete set of points. We call a point $p\in U$ critical if $f$ is not a local homeomorphism in any neighborhood of $p$. The set of critical points of $f$ will be denoted by ${\rm crit}(f)$.
We will also need the following quasiregular version of Montel’s theorem [AIM, Corollary 5.5.7] as well as a compactness property of quasisymmetric maps below.
Theorem 2.2 ([AIM]).
Suppose $z_{0},z_{1}\in\mathbb{C}$ and let $h_{k}\colon\Omega\to\mathbb{C},k=1,2,\dots$, be a sequence of $K$-quasiregular mappings defined on a domain $\Omega\subset\mathbb{C}$, each omitting these two values. Then there is a subsequence converging locally uniformly on $\Omega$ to a mapping $h$, and $h$ is either a $K$-quasiregular mapping or a constant.
Finally, recall that if $(X,d_{X})$ and $(Y,d_{Y})$ are two metric spaces, a homeomorphism $f\colon X\to Y$ is said to be $\eta$-quasisymmetric if there is an increasing homeomorphism $\eta\colon[0,\infty)\to[0,\infty)$ such that for any triple $x,y,z$ of distinct points in $X$, we have
${\frac{d_{Y}(f(x),f(y))}{d_{Y}(f(x),f(z))}}\leq\eta\left({\frac{d_{X}(x,y)}{d_{X}(x,z)}}\right).$
3. Schottky maps and their local rigidity
A relative Schottky set $S$ in a domain $D\subseteq\widehat{\mathbb{C}}$ is a set of the form
[…]
where each $B_{i},\,i\in I$, is an open geometric disk with the closure $\overline{B_{i}}$ contained in $D$, and $\overline{B_{i}}\cap\overline{B_{j}}=\emptyset$ for $i\neq j$. The boundary $\partial B_{i}$ of a disk $B_{i},\,i\in I$, is called a peripheral circle. If $D=\mathbb{C}$ or $\widehat{\mathbb{C}}$, $S$ is called a Schottky set. The study of Schottky sets was initiated in [BKM], and relative Schottky sets and maps in [Me1, Me2].
Let $S\subseteq\mathbb{C}$ be an arbitrary subset that has an accumulation point $p\in S$, and let $U$ be a neighborhood of $p$ in $\mathbb{C}$. Suppose that $f\colon S\cap U\to\mathbb{C}$ is a continuous function. We say that $f$ is conformal at $p$ if
(3) $\lim_{q\in S,\,q\to p}\frac{f(q)-f(p)}{q-p}$
exists and is non-zero. The limit above is called the derivative of $f$ at $p$ and is denoted by $f^{\prime}(p)$.
[…]
The notion of Schottky maps is closely related to that of quasiconformal maps. In fact, we will use the following criterion [Me3, Lemma 2.2] for maps to be Schottky maps.
Lemma 3.1 ([Me3]).
Let $S\subseteq\widehat{\mathbb{C}}$ be a Schottky set of measure zero. Suppose $U\subseteq\widehat{\mathbb{C}}$ is open and $f\colon U\rightarrow\widehat{\mathbb{C}}$ is a quasiconformal map with $f^{-1}(S)=S\cap U$. Then $f\colon S\cap U\to S$ is a Schottky map. In particular, if $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a quasiregular map with $f^{-1}(S)=S$, then $f\colon S\setminus\mathrm{crit}(f)\to S$ is a Schottky map.
An important property of Schottky maps is their localrigiditylocal\;rigidity, i.e., they are uniquely determined by their local behavior near one point. We need the following notion to state related results, Theorem 3.2 [Me2, Corollary 4.2] and Theorem 3.3 [Me2, Theorem 5.2] below. A relative Schottky set $S=D\setminus\cup_{i\in I}B_{i}$ is called locally porous at $p\in S$ if there exist a neighborhood $U$ of $p$ and constants $r_{0}>0,\,C\geq 1$, such that for every $q\in S\cap U$ and each $r$ with $0<r\leq r_{0}$, there exists $i\in I$ with $B(q,r)\cap B_{i}\neq\emptyset$ and
$r/C\leq{\rm diam}(B_{i})\leq Cr,$
where $B(q,r)$ denotes the open disk of radius $r$ centered at $q$, and ${\rm diam}(X)$ denotes the diameter of a set $X\subseteq\mathbb{C}$. A relative Schottky set $S$ is called locally porous if it is locally porous at every $p\in S$. Every locally porous relative Schottky set has measure zero since it cannot have Lebesgue density points.
Theorem 3.2 ([Me2]).
Let $S$ be a locally porous relative Schottky set in $D\subseteq\mathbb{C}$, and suppose that $U\subseteq D$ is an open set such that $S\cap U$ is connected. Let $f$ and $g\colon S\cap U\to\widetilde{S}$ be Schottky maps into a relative Schottky set $\widetilde{S}$ in a domain $\widetilde{D}$, and consider
$E=\{p\in S\cap U\colon f(p)=g(p)\}.$
Then $E=S\cap U$, provided $E$ has an accumulation point in $U$.
We also record another theorem [Me2, Theorem 5.2] on the rigidity of Schottky maps. Roughly speaking, it says that, under a mild assumption that is satisfied for any subhyperbolic dynamical system, a sequence $f_{k}$ of locally uniformly converging Schottky maps will become constant when $k$ is large enough.
Theorem 3.3 ([Me2]).
Let $S$ be a locally porous relative Schottky set in a domain $D\subseteq\mathbb{C}$, and $p\in S$ be an arbitrary point. Suppose that $U\subseteq D$ is an open neighborhood of $p$ such that $S\cap U$ is connected. We assume that there exists a Schottky map $g\colon S\cap U\to S$ with $g(p)=p$ and $g^{\prime}(p)\neq 1$. Let $\widetilde{S}$ be a relative Schottky set in a domain $\widetilde{D}$ and let $(f_{k})_{k\in\mathbb{N}}$ be a sequence of Schottky maps $f_{k}\colon S\cap U\to\widetilde{S}$. We assume that for each $k\in\mathbb{N}$ there exists an open set $\widetilde{U}_{k}$ so that the map $f_{k}\colon S\cap U\to\widetilde{S}\cap\widetilde{U}_{k}$ is a homeomorphism, and the sequence $(f_{k})$ converges locally uniformly to a homeomorphism $f\colon S\cap U\to\widetilde{S}\cap\widetilde{U}$, where $\widetilde{U}$ is an open set. Then there exists $N\in\mathbb{N}$ such that $f_{k}=f$ in $S\cap U$ for all $k\geq N$.
4. Conformal elevator
In [BLM, Section 4], the authors introduced the technique of conformal elevator for subhyperbolic rational maps. Here we discuss this technique for postcritically finite maps only, and we will apply it to the proof of Proposition 5.1 in Section 5.
Assume that $f$ is a postcritically finite rational map with
$\mathcal{J}(f)\subseteq\frac{1}{2}\mathbb{D}\quad and\quad f^{-1}(\mathbb{D})\subseteq\mathbb{D}.$
For a map $f$, say with an attracting fixed point, these are easy to achieve by conjugating $f$ with an appropriate Möbius transformation. In this subsection we use Euclidean metric on $\mathbb{C}$. Fix $\epsilon_{0}>0$ so small that $\mathrm{diam}(\mathcal{J}(f))>2\epsilon_{0}$, and so that every disk $B^{\prime}=B(q,r^{\prime})$ centered at a point $q\in\mathcal{J}(f)$ with radius $r^{\prime}\leq 8\epsilon_{0}$ is contained in $\mathbb{D}$, and contains at most one point in the postcritical set of $f$, i.e., in $\mathrm{post}(f)=\cup_{n\in\mathbb{N}}f^{n}({\rm crit}(f))$. Let $B=B(p,r)$ be a small disk centered at a point $p\in\mathcal{J}(f)$ and of radius $r<\epsilon_{0}$. We know that $\mathcal{J}(f)\subseteq f^{n}(B)$ for sufficiently large $n$ (see [Be, Theorem 4.2.5 (ii)]), thus $\mathrm{diam}(f^{n}(B))>2\epsilon_{0}$. In other words, there exists a maximal number $n\in\mathbb{N}$ such that $\mathrm{diam}(f^{n}(B))\leq 2\epsilon_{0}$. By uniform continuity of $f$ near the compact set $\mathcal{J}(f)$, there exists $\delta_{0}>0$ only depending on $f$ such that
$\mathrm{diam}(f^{n}(B))\geq\delta_{0}.$
Moreover, we have a lower bound for diameters of image sets $f^{n}(A)$, where $A\subseteq B$ is connected ([BLM, Lemma 4.1]):
Lemma 4.1 ([BLM]).
There exist constants $\gamma>0$ and $C>0$ only depending on $f$ and independent of $B=B(p,r)$, such that for any connected set $A\subseteq B$,
$\mathrm{diam}(f^{n}(A))^{\gamma}\geq C\frac{\mathrm{diam}(A)}{\mathrm{diam}(B)}.$
Proof.
We sketch the proof here. First, by our choice of $\epsilon_{0}$, we find a disk $B^{\prime}=B(q,r^{\prime})$ with $q\in\mathcal{J}(f)$ and $r^{\prime}\leq 8\epsilon_{0}$, such that
$f^{n}(B)\subseteq\frac{1}{2}B^{\prime}\subseteq\mathbb{D}\quad{\rm and}\quad B^{\prime}\cap\mathrm{post}(f)\subseteq\{q\}.$
Let $\Omega\in\widehat{\mathbb{C}}$ be the unique component of $f^{-n}(B^{\prime})$ that contains $B$.
Second, there exist conformal maps $\varphi\colon B^{\prime}\to\mathbb{D}$ and $\psi\colon\Omega\to\mathbb{D}$ such that
$(\varphi\circ f^{n}\circ\psi^{-1})(z)=z^{k}$
for all $z\in\mathbb{D}$, where $k\leq N=N(f)$ is uniformly bounded only by $f$ and independent of $B$. Third, by Koebe distortion theorem we have
$|u-v|\approx\left\lvert(\psi^{-1})^{\prime}(0)\right\rvert\cdot\left\lvert\varphi(u)-\varphi(v)\right\rvert$,
whenever $u,v\in B$, and where we use $a\approx b$ to indicate that there is a constant $C\geq 1$ independent of the disk $B$ such that $a/C\leq b\leq Cb$. In particular,
$\mathrm{diam}(B)\approx\left\lvert(\psi^{-1})^{\prime}(0)\right\rvert\cdot\mathrm{diam}(f^{n}(B)).$
On the other hand,
$1\approx\mathrm{diam}(f^{n}(B))$ $\approx\mathrm{diam}(\varphi(f^{n}(B)))$
$=\mathrm{diam}(P_{k}(\psi(B)))\leq C\cdot\mathrm{diam}(\psi(B))\leq 2,$
where $C>0$ is a constant independent of $B$. Hence $\mathrm{diam}(\psi(B))\approx 1$, and so $\mathrm{diam}(B)\approx\left\lvert(\psi^{-1})^{\prime}(0)\right\rvert$. This implies that
$\frac{|u-v|}{\mathrm{diam}(B)}\approx\left\lvert\varphi(u)-\varphi(v)\right\rvert$,
whenever $u,v\in B$. Hence
$\frac{\mathrm{diam}(A)}{\mathrm{diam}(B)}\approx\mathrm{diam}(\psi(A))$ $\leq C\cdot\mathrm{diam}(P_{k}(\psi(A)))^{1/N}$
$=\mathrm{diam}(\varphi(f^{n}(A)))^{1/N}\approx\mathrm{diam}(f^{n}(A))^{1/N},$
for another constant $C>0$ independent of $B$.
5. Local quasiconformal maps between carpet Julia sets
In this section, $f$ and $g$ are postcritically finite rational maps, and the corresponding Julia sets $\mathcal{J}(f)$ and $\mathcal{J}(g)$ are Sierpiński carpets. In [BLM, Theorem 1.4], the authors proved that a quasisymmetric map $\xi\colon\mathcal{J}(f)\to\mathcal{J}(g)$ is the restriction to $\mathcal{J}(f)$ of a Möbius transformation. Many of the arguments in that paper are local and it makes sense to investigate whether they carry over to quasiregular maps $\xi$ that restrict to maps between $\mathcal{J}(f)$ and $\mathcal{J}(g)$. Note that a quasisymmetric map $\xi\colon\mathcal{J}(f)\to\mathcal{J}(g)$ extends to a global quasiconformal map of $\widehat{\mathbb{C}}$. One of the main ingredients of [BLM] is the functional equation [BLM, Equation (1.2)]
$g^{m^{\prime}}\circ\xi=g^{m}\circ\xi\circ f^{n},$
established on all of $\mathcal{J}(f)$ in [BLM, Section 8, Step IV]. Since quasiregular maps are locally quasiconformal away from critical points, we obtain such a functional equation in the local setting; see Proposition 5.1 below, and then conclude that it holds globally on $\mathcal{J}(f)$ for quasiregular maps in Theorem 5.2. We outline the proofs with these modifications for readers’ convenience.
Proposition 5.1.
Let $f,g\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be postcritically finite rational maps, and suppose that the corresponding Julia sets $\mathcal{J}(f)$ and $\mathcal{J}(g)$ are Sierpiński carpets. If $U$ and $V$ are open sets non-trivially intersecting $\mathcal{J}(f)$ and $\mathcal{J}(g)$, respectively, and $\xi\colon U\rightarrow V$ is a quasiconformal map with $\xi^{-1}(\mathcal{J}(g)\cap V)=\mathcal{J}(f)\cap U$, then there exist an open set $W\subseteq U$ with $\mathcal{J}(f)\cap W\neq\emptyset$, and $n,m,m^{\prime}\in\mathbb{N}$, such that on $\mathcal{J}(f)\cap W$ one has
(4) $g^{m^{\prime}}\circ\xi=g^{m}\circ\xi\circ f^{n}.$
Proof.
By conjugating $g$ (hence $\mathcal{J}(g)$), we may assume that $\mathcal{J}(g)\subset\mathbb{D}$. Below we use the Euclidean metric. Since repelling periodic points of $f$ are dense in $\mathcal{J}(f)$ (see [Be, p. 148, Theorem 6.9.2]), we can find a repelling periodic point $p\in\mathcal{J}(f)\cap U$ of $f$ of period $n$ and $\rho>0$ such that $U_{0}=B(p,\rho)\subseteq U$ is disjoint from $\mathrm{post}(f)$. Denote $n\cdot k$ by $n_{k}$, and the connected component of $f^{-n_{k}}(U_{0})$ containing $p$ by $U_{k}$. By possibly taking a smaller $\rho$, we may assume that $U_{0}\supseteq U_{1}\supseteq U_{2}\supseteq$... and $\mathrm{diam}(U_{k})\to 0$ as $k\to\infty$. We take the branch $f^{-n_{k}}$ on $U_{0}$ to be the inverse map of $f^{n_{k}}|_{U_{k}}\colon U_{k}\to U_{0}$. Then the choice of these inverse branches is consistent in the sense that we have
(5) $f^{n_{k+1}-n_{k}}\circ f^{-n_{k+1}}=f^{-n_{k}}$
for all $k$. Let $r_{k}>0$ be the smallest number such that $\xi(U_{k})\subseteq B(\xi(p),r_{k})$. Since $\xi(p)\in\xi(U_{k})$, $\mathrm{diam}(\xi(U_{k}))\approx r_{k}$, where, similar to above, the symbol $\approx$ between two quantities means that their quotient is bounded above and below by a constant independent of $k$. By applying Lemma 4.1 to $A=\xi(U_{k})$ and $B=B(\xi(p),r_{k})$, we can find iterates $g^{m_{k}}$ such that $\mathrm{diam}\left(g^{m_{k}}(\xi(U_{k}))\right)\approx 1$. Since $g^{m_{k}}(\xi(U_{k}))$ meets $\mathcal{J}(g)$ at $g^{m_{k}}(\xi(p))$, it is contained in a bounded region of $\mathbb{C}$.
Now we consider the composition
$h_{k}=g^{m_{k}}\circ\xi\circ f^{-n_{k}}$
defined on $B(p,\rho)$ for $k\in\mathbb{N}$. Since $h_{k}$ are uniformly $K$-quasiregular ($K$ only depends on $\xi$) maps from $U_{0}$ into a compact subset of $\mathbb{C}$, Theorem 2.2 implies that a subsequence of $\{h_{k}\}$ subconverges locally uniformly to a map $h\colon U_{0}\to\mathbb{C}$ that is also quasiregular or a constant. The latter possibility is ruled out since $\mathrm{diam}\left(h_{k}(U_{0})\right)=g^{m_{k}}(\xi(U_{k}))\approx 1$. For simplicity of notations we assume that the whole sequence $h_{k}$ converges; though $n_{k}$ is no longer $n\cdot k$ in this case. The map $h$ has at most countably many critical points, and so there exists a non-critical point $q\in\mathcal{J}(f)\cap U_{0}$ and a small neighborhood $B(q,r)$ of $q$ such that $h$ is injective on $B(q,r)$, and hence quasiconformal. In what follows, we replace $U_{0}$ by this $B(q,r)$.
We know by [BLM, Theorem 1.10] that the peripheral circles of a Sierpiński carpet Julia set $\mathcal{J}(f)$ of a postcritically finite rational map $f$ satisfy nice geometric properties. Namely, they are uniformly relatively separated uniform quasicircles and appear at all locations and scales; see [BLM] for definitions. In particular, according to [Bo, Corollary 1.2], there exists a quasisymmetric map $\beta$ on $\widehat{\mathbb{C}}$ such that $S=\beta(J(f))$ is a round Sierpiński carpet, i.e., its peripheral circles are geometric circles. Moreover, $S$ is locally porous.
From Lemma 3.1 we know that any iterate
$f_{\beta}^{n}=\beta\circ f^{n}\circ\beta^{-1}\colon S\backslash\mathrm{crit}(f_{\beta}^{n})\to S$
is a Schottky map. Moreover, since
$h^{-1}(\mathcal{J}(g))=h_{k}^{-1}(\mathcal{J}(g))=\mathcal{J}(f)\cap U_{0},$
by the same lemma, the quasiconformal maps $\beta\circ h\circ\beta^{-1},\beta\circ h_{k}\circ\beta^{-1},\ k\in\mathbb{N}$, are Schottky maps on $S\cap\beta(U_{0})$.
It is easy to see ([Me3, Lemma 2.1]) that one can find a small Jordan region $W^{\prime}$ with $a\in W^{\prime}\subseteq\beta(U_{0})$ such that $S\cap W^{\prime}$ is connected. Since repelling periodic points are dense in $\mathcal{J}(f)$, we can find a point $a\in\left(S\backslash\mathrm{crit}(f_{\beta}^{n})\right)\cap W^{\prime}$ and $n\in\mathbb{N}$ such that
$f_{\beta}^{n}(a)=a\quad\mathrm{and}\quad(f_{\beta}^{n})^{\prime}(a)\neq 1.$
Theorem 3.3 implies that
$\beta\circ h_{k}\circ\beta^{-1}\equiv\beta\circ h\circ\beta^{-1}\quad\mathrm{on}\quad S\cap W^{\prime},$
for $k$ large enough. Hence $h_{k}=h_{l}$ on $\mathcal{J}(f)\cap W^{\prime\prime}$, where $W^{\prime\prime}=\beta^{-1}(W^{\prime})$, for some $k<l$, i.e.,
$g^{m_{k}}\circ\xi\circ f^{-n_{k}}=g^{m_{l}}\circ\xi\circ f^{-n_{l}}\quad\mathrm{on}\quad\mathcal{J}(f)\cap W^{\prime\prime}.$
Finally, we pre-compose $f^{n_{l}}$ on both sides and from Equation (5) get
$g^{m_{k}}\circ\xi\circ f^{n_{l}-n_{k}}=g^{m_{l}}\circ\xi\quad\mathrm{on}\quad\mathcal{J}(f)\cap W,$
where $W=f^{-n_{l}}(W^{\prime\prime})$, establishing (4).
As a consequence we obtain the following result.
Theorem 5.2.
If $f$ and $g$ are postcritically finite rational maps whose respective Julia sets $\mathcal{J}(f),\mathcal{J}(g)$ are Sierpiński carpets, and if $\xi$ is a quasiregular map $\xi\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ such that $\xi^{-1}(\mathcal{J}(g))=\mathcal{J}(f)$, then Equation (4) holds everywhere on $\mathcal{J}(f)$.
Proof.
Since $\mathcal{J}(g)$ has the same properties as $\mathcal{J}(f)$, by [Bo, Corollary 1.2] there exists a quasisymmetric map $\gamma$ on $\widehat{\mathbb{C}}$ such that $\widetilde{S}=\beta(J(g))$ is a round Sierpiński carpet. Using the same notations and arguments as in Proposition 5.1 and its proof, we know that $H_{1}=\gamma\circ g^{m^{\prime}}\circ\xi\circ\beta^{-1}$ and $H_{2}=\gamma\circ g^{m}\circ\xi\circ f^{n}\circ\beta^{-1}$ are Schottky maps from $S\backslash(\mathrm{crit}(H_{1})\cup\mathrm{crit}(H_{2}))$ to $\widetilde{S}$. Proposition 5.1 then says that the set
$E=\left\{p\in S\backslash(\mathrm{crit}(H_{1})\cup\mathrm{crit}(H_{2}))\;|\;H_{1}(p)=H_{2}(p)\right\}$
has an accumulation point in $S\backslash(\mathrm{crit}(H_{1})\cup\mathrm{crit}(H_{2}))$. According to Theorem 3.2, we have $H_{1}=H_{2}$ on $S\backslash(\mathrm{crit}(H_{1})\cup\mathrm{crit}(H_{2}))$. By continuity it holds everywhere on $S$, and the conclusion follows.
6. Rational maps sharing a measure of maximal entropy
It is an interesting problem to study rational maps that share a non-trivial Julia set, or have the same measure of maximal entropy. Here we briefly summarize relevant notions and basic properties. The reader can refer to [DS] for details.
If $f$ is a rational map of $\widehat{\mathbb{C}}$ of degree $d$, then the measure of maximal entropy $\mu_{f}$ of $f$ is the unique pull-back invariant probability measure with support equal to the Julia set $\mathcal{J}(f)$ of $f$, achieving maximal entropy $\mathrm{log}\,d$. Pull-back invariant means that $\frac{1}{d}f^{*}\mu_{f}=\mu_{f}$, or $\mu_{f}(f(A))=d\cdot\mu_{f}(A)$, whenever $f$ is injective on a Borel set $A$. Note that this implies a weaker notion of invariance that $f_{*}\mu_{f}=\mu_{f}$, or $\mu_{f}(f^{-1}(A))=\mu_{f}(A)$ for all Borel sets $A\subseteq\widehat{\mathbb{C}}$.
For any probability measure $\mu$ on $\widehat{\mathbb{C}}$ and a rational map $f$, the Jacobian $J_{\mu}(f)$ of $f$ is defined as the Radon-Nikodym derivative of $f^{*}(\mu)$ with respect to $\mu$. Namely, for each Borel measurable set $A$ such that $f$ is injective on $A$,
$\mu(f(A))=\int_{A}J_{\mu}(f)(x)\>\mathrm{d}\mu(x),\quad\mathrm{whenever}\;f|_{A}\;\mathrm{is}\;\mathrm{injective}.$
If $\mu=\mu_{f}$ is the measure of maximal entropy of a rational map $f$ of degree $\mathrm{deg}(f)$ at least 2, then $J_{\mu_{f}}(f)$ is constant equal to $\mathrm{deg}(f)$. In fact, the converse is generally true ([DS, Theorem 1.118]):
Theorem 6.1 ([DS]).
Let $f$ be a ratinal map of degree $d$. If $\mu$ is a pull-back invariant probability measure of constant Jacobian $d$, then $\mu=\mu_{f}$.
Proof.
The measure $\mu_{f}$ is the unique probability measure such that the entropy $h_{\mu}(f)$ attains the maximum value $\mathrm{log}\,d$. The result now follows from the inequality [Pa]
$h_{\nu}(f)\geq\int\mathrm{log}J_{\nu}(f)\;\mathrm{d}\nu,$
satisfied for any probability measure $\nu$ invariant under $f$, i.e., such that $f_{*}\nu=\nu$. Indeed, by setting $\nu=\mu$ from the statement of the theorem we get $h_{\mu}(f)\geq\mathrm{log}\,d$, and so $h_{\mu}(f)$ attains the maximum. Therefore $\mu=\mu_{f}$.
It is well known that $\mu_{f}=\mu_{f^{n}}$ for all $n\in\mathbb{N}$, and commuting rational maps have the same measure of maximal entropy. The following theorem is contained in [Le, Le2, LP]; see also [Ye, Theorem 1.7].
Theorem 6.2 ([Le, Le2, LP, Ye]).
Let $f,g\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be non-exceptional rational maps of degrees at least 2. Then $\mu_{f}=\mu_{g}$ if and only if for some $k,l\in\mathbb{N}$,
$f^{k}\circ g^{l}=f^{2k}.$
Here, a rational map is called exceptional if it is conformally conjugate to either a power map, or a Chebyshev polynomial, or a Lattès map. In particular, a rational map whose Julia set is a Sierpiński carpet is non-exceptional.
For general rational maps $f$ and $g$ sharing the same Julia set, they need not have common iterates up to Möbius transformations ([Ye, Theorem 1.1]). In general, holomorphic maps of the form $f^{-k}\circ f^{k}$ need not be Möbius, as the following example shows.
Example.
Consider $f(z)=z^{2}-\frac{1}{16z^{2}}$. It is postcritically finite and has a Sierpiński carpet Julia set. A priori multi-valued map $f^{-1}\circ f(z)$ must be either $\pm z$ or $\pm\frac{i}{4z}$, i.e., a Möbius transformation. However, when $k\geq 2$, $f^{-k}\circ f^{k}$ first defined locally and then extended analytically may not be a Möbius transformation. For example, by solving for the coefficients of a Möbius transformation $\xi$ such that
(6) $f^{2}\circ\xi=f^{2},$
one can show that $\xi$ belongs to the group generated by $\rho(z)=\mathrm{i}z$ and $\iota(z)=\frac{1}{4z}$. Indeed, if $\xi(z)={(az+b)}/{(cz+d)}$ satisfies (6), we may assume $a\neq 0$, otherwise we may consider $\iota\circ\xi$. Then
$\lim_{z\to\infty}f^{2}(\xi(z))=\lim_{z\to\infty}f^{2}(z)=\infty,$
while $\lim_{z\to\infty}\xi(z)=\frac{a}{c}$. Hence $f^{2}(\frac{a}{c})=\infty$, implying $c=0$. We may assume $d=1$, and thus $\xi(z)=az+b$. But $f^{2}(z)=z^{4}+O(1)$ and so $f^{2}(az+b)\thicksim(az+b)^{4}+O(1)$ as $z\to\infty$. These two functions being identical, it implies that $a^{4}=1$ and $b=0$ and the claim follows. In particular, the group of Möbius transformation satisfuing (6) has eight elements. On the other hand, the map $f^{-2}$ has 16 branches in a neighborhood of a noncritical value for $f^{2}$, so there must be some local holomorphic maps of the form $f^{-2}\circ f^{2}$ which are not Möbius.
7. Proofs of Theorems 1.2 and 1.3
We need the following lemma in the proof of Theorem 1.2; cf. [BLM, Lemma 7.1].
Lemma 7.1.
Let $\phi\colon\partial\mathbb{D}\rightarrow\partial\mathbb{D}$ be an orientation preserving branched covering, and suppose that there exist numbers $k,l,n\in\mathbb{N}$, $k\geq 2$, such that
(7) $(P_{l}\circ\phi)(z)=(P_{n}\circ\phi\circ P_{k})(z)\quad\text{for $z\in\partial\mathbb{D}$},$
where $P_{n}(z)=z^{n}$ for all $n\in\mathbb{N}$. Then $l=n\cdot k$ and there exists $b\in\partial\mathbb{D},m\in\mathbb{N}$, such that $\phi(z)=bz^{m}$ for all $z\in\partial\mathbb{D}$.
Proof.
By counting topological degrees of both sides, one immediately sees that $l=n\cdot k$. Let $m$ be the degree of $\phi$. Then there exists a continuous function $\alpha\colon\mathbb{R}\rightarrow\mathbb{R}$ with $\alpha(t+2\pi)=\alpha(t)+2\pi m$ such that
$\phi(\mathrm{e}^{\mathrm{i}t})=\mathrm{e}^{\mathrm{i}\alpha(t)}\quad\text{for $t\in\mathbb{R}$}.$
By (7) we have
$\mathrm{e}^{\mathrm{i}l\alpha(t)}=\mathrm{e}^{\mathrm{i}n\alpha(kt)}$
for $t\in\mathbb{R}$. This implies that there exists a constant $c\in\mathbb{R}$ such that
$\alpha(t)=\frac{n}{l}\alpha(kt)+c=\frac{1}{k}\alpha(kt)+c\quad\text{for $t\in\mathbb{R}$}.$
Using induction on $s\in\mathbb{N}\cup\{0\}$, one obtains
$\alpha\left(t+\frac{2\pi}{k^{s}}\right)=\frac{1}{k}\alpha\left(kt+\frac{2\pi}{k^{s-1}}\right)+c=\frac{1}{k}\alpha(kt)+\frac{2\pi}{k^{s}}m+c=\alpha(t)+\frac{2\pi}{k^{s}}m,$
for all $t\in\mathbb{R}$. Continuity of $\alpha$ then implies that $\alpha$ is a linear function and the conclusion follows.
7.1. Proof of Theorem 1.2
We follow the steps, namely steps I-VI, in [BLM, Section 8].
Steps I-IV are concerned with establishing equation
(8) $g^{m^{\prime}}\circ\xi=g^{m}\circ\xi\circ f^{n},$
for some $n,m$, and $m^{\prime}$, on all of $\mathcal{J}(f)$. Proposition 5.1 and Theorem 5.2 above do just that under the assumption that $\xi\colon\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ is a quasiregular map with $\xi^{-1}(\mathcal{J}(g))=\mathcal{J}(f)$.
Following step V of [BLM, Section 8], we derive Equation (1). Indeed, by post-composing both sides of Equation (8) by a suitable iterate of $g$ and denoting $f^{n}$ by $F$ and $g^{m^{\prime}-m}$ by $G$, we obtain
(9) $G^{l+1}\circ\xi=G^{l}\circ\xi\circ F$
on $\mathcal{J}(f)$ for some $l\in\mathbb{N}$. This equation immediately implies Equation (1) on $\mathcal{J}(f)$.
The final step VI of [BLM, Section 8] defines a natural extension of $\xi$ from $\mathcal{J}(f)$ to $\widehat{\mathbb{C}}$ that is analytic in each Fatou component of $\mathcal{J}(f)$. It is obtained by first extending $\xi|_{\mathcal{J}(f)}$ in each periodic Fatou component of $\mathcal{J}(f)$ and then lifting via the dynamics to the remaining Fatou components, which are necessarily pre-periodic by Sullivan’s No Wandering Domains Theorem. In [BLM] such extension is actually conformal in each Fatou component by [BLM, Lemma 7.1]. In our setting, i.e., when $\xi$ is assumed to be quasiregular, we obtain only an analytic extension to the periodic Fatou components, in a similar way using Böttcher coordinates and Lemma 7.1 in place of [BLM, Lemma 7.1].
Indeed, first let $U$ be a periodic Fatou component of $F$ from (9). We denote by $k\in\mathbb{N}$ the period of $U$, and define $V$ to be the Fatou component of $G$ bounded by $\xi(\partial U)$, and $W=G^{l}(V)$, where $G$ and $l\in\mathbb{N}$ are as in (9). Then (9) implies that $W$ is invariant under $G^{k}$. By [BLM, Lemma 3.3], the basepoint-preserving homeomorphisms $\psi_{U}\colon(\overline{U},p_{U})\rightarrow(\overline{\mathbb{D}},0)$ and $\psi_{W}\colon(\overline{W},p_{W})\rightarrow(\overline{\mathbb{D}},0)$, for unique $p_{U}\in U$ and $p_{W}\in W$, conjugate $F^{k}$ and $G^{k}$, respectively, to power maps. Since $f$ and $g$ are postcritically finite, the periodic Fatou components $U$ and $W$ are superattracting, and thus the degrees of these power maps are at least 2.
Again by Lemma  [BLM, Lemma 3.3], there exists $\psi_{V}\colon(\overline{V},p_{V})\to(\overline{\mathbb{D}},0),\ p_{V}\in V$, such that $\psi_{W}\circ G^{l}\circ\psi^{-1}_{V}$ is a power map. Since $U,V,W$ are Jordan domains, the maps $\psi_{U},\psi_{V},\psi_{W}$ give homeomorphisms between the boundaries of the corresponding Fatou components and $\partial\mathbb{D}$. Since $\xi$ is a branched covering of $\partial U$ onto $\partial V$, the map $\phi=\psi_{V}\circ\xi\circ\psi_{U}^{-1}$ gives a branched covering on $\partial\mathbb{D}$. Now (9) implies that on $\partial\mathbb{D}$ we have
$P_{d_{3}}\circ\phi$ $=\psi_{W}\circ G^{k+l}\circ\psi_{V}^{-1}\circ\phi=\psi_{W}\circ G^{k+l}\circ\xi\circ\psi_{U}^{-1}$
$=\psi_{W}\circ G^{l}\circ\xi\circ F^{k}\circ\psi^{-1}_{U}=\psi_{W}\circ G^{l}\circ\xi\circ\psi^{-1}_{U}\circ P_{d_{1}}$
$=\psi_{W}\circ G^{l}\circ\psi^{-1}_{V}\circ\phi\circ P_{d_{1}}=P_{d_{2}}\circ\phi\circ P_{d_{1}},$
for some $d_{1},d_{2},d_{3}\in\mathbb{N}$ with $d_{1}\geq 2$. Lemma 7.1 implies that $\phi$ extends to $\overline{\mathbb{D}}$ as a monomial around 0, also denoted by $\phi$. In particular, $\phi(0)=0$, and so $\phi$ preserves the basepoint 0 in $\overline{\mathbb{D}}$. If we define $\xi=\psi_{V}^{-1}\circ\phi\circ\psi_{U}$ on $\overline{U}$, then $\xi$ is a branched covering of $(\overline{U},p_{U})$ onto $(\overline{V},p_{V})$.
In this way, we can analytically extend $\xi$ to every periodic Fatou component of $F$ so that $\xi$ maps the basepoint of a Fatou component to the basepoint of the image component. To get an extension also for other, non-periodic, Fatou components of $F$, we proceed inductively on the level of such a given Fatou component $V$, i.e., the smallest $n\in\mathbb{N}\cup\{0\}$, such that $F^{n}(V)$ is periodic. Suppose that the level of $V$ is $n\in\mathbb{N}$ and that we already have an extension for all Fatou components with level smaller than $n$. This applies to the Fatou component $U=F(V)$ of $F$, and so a conformal extension $(U,p_{U})\rightarrow(U^{\prime},p_{U^{\prime}})$ of $\xi|_{\partial U}$ exists, where $U^{\prime}$ is the Fatou component of $G$ bounded by $\xi(\partial U)$. Let $V^{\prime}$ be the Fatou component of $G$ bounded by $\xi(\partial V)$, and $W=G^{l+1}(V^{\prime})$. Then, by using (9) on $\partial V$ we conclude that $G^{l}(U^{\prime})=W$. Define $\alpha=G^{l}\circ\xi|_{\overline{U}}\circ F|_{\overline{V}}$ and $\beta=\xi|_{\partial V}$. Then the assumptions of the Lifting Lemma [BLM, Lemma 3.4] are satisfied for $D=V$, $p_{D}=p_{V}$, and the iterate $G^{l+1}\colon V^{\prime}\rightarrow W$ of $G$. Indeed, $\alpha$ is continuous on $\overline{V}$ and holomorphic on $V$, we have
$\alpha^{-1}(p_{W})=F|_{\overline{V}}^{-1}(\xi|_{\overline{U}}^{-1}(p_{U^{\prime}}))=F|_{\overline{V}}^{-1}(p_{U})=p_{V},$
and
$G^{l+1}\circ\beta=G^{l+1}\circ\xi|_{\partial V}=G^{l}\circ\xi|_{\partial U}\circ F|_{\partial V}=\alpha|_{\partial V}.$
Since $\beta$ is continuous, [BLM, Lemma 3.4] implies that there exists a continuous map $\widetilde{\alpha}$ of $(\overline{V},p_{V})$ onto $(\overline{V}^{\prime},p_{V^{\prime}})$ such that $\widetilde{\alpha}|_{\partial V}=\beta=\xi|_{\partial V}$. In other words, $\widetilde{\alpha}$ gives the desired basepoint preserving analytic extension to the Fatou component $V$.
[…]
Finally, if we replace $\xi$ by $\widehat{\xi}$, both sides of Equation (9) are rational maps that agree on $\mathcal{J}(f)$. Therefore they agree on all of $\widehat{\mathbb{C}}$, establishing Equation (1). ∎
7.2. Proof of Theorem 1.3
[…]
$J_{\mu}(\widehat{\xi})(x)=J_{\mu}(f^{k}\circ\widehat{\xi}(x))/J_{\mu}(f^{k})(\widehat{\xi}(x))=\mathrm{const}/\mathrm{const}=\mathrm{const},$
i.e., $(\widehat{\xi})^{*}(\mu)=\mathrm{deg}(\widehat{\xi})\mu$. By Theorem 6.1, this implies $\mu=\mu_{\widehat{\xi}}$, and in particular $\mathcal{J}(\widehat{\xi})=\mathcal{J}(f^{k})=\mathcal{J}(f)$. Theorem 6.2 applied to $f$ and $\widehat{\xi}$ gives Equation (2). Equation (2) in turn gives that the critical points of $\widehat{\xi}$ are among the critical points of an iterate of $f$, and indeed $\widehat{\xi}$ is postcritically finite. ∎
Several comments are in order. If $\xi$ is a quasiconformal map of $\widehat{\mathbb{C}}$ in Theorem 1.3, then [BLM, Theorem 1.1] gives that $\xi$ is the restriction of a Möbius transformation $\widehat{\xi}$ to $\mathcal{J}(f)$. From [Le, Theorem 1.1] we also know that then $\widehat{\xi}^{m}=\mathrm{id}$ for some $m\in\mathbb{N}$, and the arguments from [BLM] give that $\widehat{\xi}$ satisfies the functional equation
$f^{2k}\circ\widehat{\xi}=f^{k}\circ\widehat{\xi}\circ f^{k}\;\mathrm{for\ some}\;k\in\mathbb{N};$
see also the proof of Theorem 1.3 above. Replacing $f$ by its iterate, we can assume $f^{2}\circ\widehat{\xi}=f\circ\widehat{\xi}\circ f$.
Question 7.1.
Let $f$ be a postcritically finite rational maps whose Julia set $\mathcal{J}(f)$ is a Siepiński carpet, and let $\widehat{\xi}$ be a Möbius transformation preserving $\mathcal{J}(f)$, and such that $f^{2}\circ\widehat{\xi}=f\circ\widehat{\xi}\circ f$. Does it imply that $F\circ\widehat{\xi}=F$, where $F$ is an iterate of $f$?
In fact, it is unclear whether such relations hold for specific maps such as $f(z)=z^{2}-\frac{1}{16z^{2}}$. However, we suspect that the following question has a positive answer.
Question 7.2.
Let $f(z)=z^{2}-\frac{1}{16z^{2}}$, and $\rho,\iota$ be as in the example of Section 6, i.e., $\rho(z)=\mathrm{i}z$ and $\iota(z)=\frac{1}{4z}$. Is it true that the group of all Möbius transformations preserving $\mathcal{J}(f)$ is $\left\langle\rho,\iota\right\rangle$?
8. Quasiregular maps between other topological types of Julia sets
One can ask whether similar statements to those in Theorems 1.2 and Theorem 1.3 hold for other topological types of non-exceptional postcritically finite rational maps. Here, again, by exceptional maps we mean, up to conjugation, power maps, Chebyshev polynomials, or Lattès maps. In this section we discuss two examples: the basilica and the Apollonian gasket cases. In contrast to the Siepiński carpet case, we demonstrate that a quasiregular map of the basilica or the gasket Julia sets need not be related to the dynamics of the corresponding map, contrasting the Sierpiński carpet case of Theorems 1.2 and 1.3.
Example.
The basilica; see [LM] for a description of the quasisymmetry group of the basilica and the discussion of the Thompson group $T$, and [BF] for more general statements and results. The basilica is the filled Julia set $\mathcal{K}(f)$ of the quadratic polynomial $f(z)=z^{2}-1$. This polynomial has a superattracting cycle $\gamma=\{0,1\}$. By definition, the immediate basin of $\gamma$ is the union of two components $U_{0}$ and $U_{-1}$ of the Fatou set containing 0 and -1, respectively. These components are Jordan disks, and they have a unique intersection point $\alpha=\frac{1-\sqrt{5}}{2}$, which is a fixed point of $f$.
Let $U_{\infty}$ denote the basin at infinity, i.e., the unbounded Fatou component of $f$, and let $\phi_{\infty}$ be the Böttcher coordinate of $U_{\infty}$. Namely, $\phi_{\infty}$ is the conformal map of $U_{\infty}$ onto $\mathbb{C}\backslash\overline{\mathbb{D}}$ conjugating $f$ to $g(z)=z^{2}$. It is known that the Julia set of a hyperbolic map $f$ is locally connected, so $\psi_{\infty}=\phi_{\infty}^{-1}\colon\mathbb{C}\backslash\overline{\mathbb{D}}\to U_{\infty}$ exteneds continuously to the boundary, and so induces a continuous boundary map $\psi$ of $\mathbb{S}^{1}$ onto the Julia set $\mathcal{J}(f)$ of $f$ satisfying $\psi\circ g=f\circ\psi$. Since $\alpha$ is a fixed point of $f$, it can be calculated that $\psi^{-1}(\alpha)=\{\mathrm{e}^{\frac{2}{3}\pi\mathrm{i}},\mathrm{e}^{\frac{4}{3}\pi\mathrm{i}}\}$.
Since the interior of $\mathcal{K}(f)$ for $f(z)=z^{2}-1$ equals the (full, not just immediate) basin of the cycle $\gamma$, every bounded Fatou component $U$ of $f$ eventually, i.e., under a certain iterate of $f$, lands in the cycle $\{U_{0},U_{-1}\}$. In fact, for each such bounded Fatou component $U\neq U_{0}$ there exists a unique $n\in\mathbb{N}$ such that $f^{n}\colon U\to U_{0}$ is a conformal map. Let $L(U)$ denote the limb of $U$, namely the closure of the union of $U$ and all the components of the interior of $\mathcal{K}(f)$ separated by $U$ from $U_{0}$. Note here that the basilica has a tree-like structure. Then $\psi^{-1}(\partial L(U))$ maps under $g^{n}$ onto ${e}^{-\frac{2}{3}\pi{i}},{e}^{\frac{2}{3}\pi{i}}]\subseteq\mathbb{S}^{1}$. The two endpoints ${e}^{\pm\frac{2}{3}\pi{i}}$ are mapped to the same point under $\psi$, called the root of $L(U)$.
In [LM, Lemma 5.4], it is shown that the Thompson group $T$ acts by quasisymmetries on $\mathcal{J}(f)$, where $f(z)=z^{2}-1$, leaving $\partial U_{0}$ invariant. In fact, such an action is defined piecewise using forward iterates of $f^{2}$ and branches of $f^{-2n},\ n\in\mathbb{N}$. Thompson group elements shuffle the limbs rooted at $\partial U_{0}$. Moreover, these quasisymmetries extend to global quasiconformal maps of $\mathbb{C}$. It is immediate that if one takes a non-trivial element of the Thompson group acting by a quasisymmetry $\xi$ on $\mathcal{J}(f)$, say that fixes all the limbs rooted on a certain non-trivial arc of $\partial U_{0}$, then this $\xi$ does not relate globally to the dynamics of $f$, say as in Equation (2). Such a map $\xi$ has degree one however. Post-composing $\xi$ with $f$ produces a quasiregular map of degree two that does not satisfy Equation (2) on $\mathcal{J}(f)$.
Example.
The Apollonian gasket; see [LLMM] for the discussion of the quasisymmetry group of the Apollonian gasket Julia set, and also [LMMN] for more general statements.
We need the following quasiregular extension lemma; see, e.g., [LN, Lemma 2.2] for similar results.
Lemma 8.1.
Let $f$ be a continuous orientation preserving branched covering map of the unit circle $\mathbb{S}^{1}$, and for $n\in\mathbb{N}$, let $I_{1},I_{2},\dots,I_{n}$, be closed arcs of $\mathbb{S}^{1}$ with disjoint interiors and whose union is $\mathbb{S}^{1}$. Assume that $f|_{I_{j}}$ extends in a neighborhood of $I_{j}$ to a holomorphic function with non-vanishing derivative for each $j=1,2,\dots,n$. Then $f$ can be extended to a quasiregular map $\widehat{f}\colon\overline{\mathbb{D}}\to\overline{\mathbb{D}}$.
Proof.
Let $e^{i\theta_{j}},\ j=1,2,\dots,n$, be the points on $\mathbb{S}^{1}$ such that $e^{i\theta_{j}}\in I_{j}\cap I_{j+1},\ j=1,2,\dots n-1$, and $e^{i\theta_{n}}\in I_{n}\cap I_{1}$. We define $\widehat{f}(z)=|z|f(\frac{z}{|z|})$ on $\overline{\mathbb{D}}\setminus\{0\}$, and $\widehat{f}(0)=0$. It is a continuous map extending $f$. If $z\in\mathbb{D}$ and ${\rm arg}(z)\notin\{\theta_{1},\theta_{2},\dots,\theta_{n}\}$, where ${\rm arg}(z)$ denotes the argument of $z$, we have
$\partial_{\bar{z}}\widehat{f}=\frac{z}{2|z|}f\left(\frac{z}{|z|}\right)-\frac{1}{2}\left(\frac{z}{|z|}\right)^{2}f^{\prime}\left(\frac{z}{|z|}\right),$
$\partial_{z}\widehat{f}=\frac{\bar{z}}{2|z|}f\left(\frac{z}{|z|}\right)+\frac{1}{2}f^{\prime}\left(\frac{z}{|z|}\right).$
Since $f$ maps $\mathbb{S}^{1}$ to itself and preserves the orientation, we get that $\mathrm{arg}(f^{\prime}(z))+\mathrm{arg}(z)\equiv\mathrm{arg}(f(z))\;\mathrm{mod}\;2\pi$ for each $z\in\mathbb{S}^{1},\ {\rm arg}(z)\notin\{\theta_{1},\theta_{2},\dots,\theta_{n}\}$. Therefore, for $z\in\mathbb{D},\ {\rm arg}(z)\notin\{\theta_{1},\theta_{2},\dots,\theta_{n}\}$,
[…]
By the assumption, there exists $\varepsilon>0$ such that $|f^{\prime}(z)|\geq\varepsilon$ for any $z\in\mathbb{S}^{1},\ {\rm arg}(z)\notin\{\theta_{1},\theta_{2},\dots,\theta_{n}\}$, and hence
$\left\lvert\frac{\partial_{\bar{z}}\widehat{f}}{\partial_{z}\widehat{f}}\right\rvert=\frac{\left\lvert 1-\left\lvert f^{\prime}\left(\frac{z}{|z|}\right)\right\rvert\right\rvert}{1+\left\lvert f^{\prime}\left(\frac{z}{|z|}\right)\right\rvert}\leq k<1,$
for some $k$ independent of $z$. From the assumptions and the extension formula it is immediate that $\widehat{f}$ is absolutely continuous on lines, and hence $\widehat{f}$ is a quasiregular extension of $f$.
Now we are able to prove the following result.
Theorem 8.2.
There exists a postcritically finite rational map $f\colon\mathbb{C}\to\mathbb{C}$ with gasket Julia set $\mathcal{J}(f)$, and a global quasiregular map $\xi\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ with $\xi^{-1}(\mathcal{J}(f))=\mathcal{J}(f)$, such that $\xi|_{\mathcal{J}(f)}$ does not have a rational extension to $\widehat{\mathbb{C}}$.
Proof.
Let $f(z)=\frac{3z^{2}}{2z^{3}+1}$ considered in [LLMM]. Its critical points are $0,1,\omega$ and $\omega^{2}$, where $\omega=e^{2\pi i/3}$, with orbits $0\mapsto 0$, $1\mapsto 1$, and $\omega\mapsto\omega^{2}\mapsto\omega$. In particular, $f$ is postcritically finite. Note that $\bar{f}$, where the bar denotes the complex conjugation, is critically fixed, i.e., all its critical points are fixed. The Julia set $\mathcal{J}(f)=\mathcal{J}(\bar{f})$ is the Apollonian gasket depicted in Figure 1.
Figure 1. The Julia set of $f(z)=\frac{3z^{2}}{2z^{3}+1}$.
Denote the central closed triangular region in Figure 1 by $\Delta ABC$, with three edge $AB,BC,CA$ being arcs on the boundaries of the disks $D_{1},D_{2},D_{3}$, respectively. The map $f^{2}$ fixes the points $A,B,C$, and $D_{1},D_{2},D_{3}$, are its fixed Fatou components. Now we take
$\xi(z)=\begin{cases}f^{2}(z),\quad z\in\Delta ABC\\ z,\quad z\in\widehat{\mathbb{C}}\backslash\left((\Delta ABC)\cup\sqcup_{i=1}^{3}D_{i}\right).\end{cases}$
The map $\xi$ can be extended to a quasiregular map of $\widehat{\mathbb{C}}$. Indeed, the boundaries of $D_{1},D_{2},D_{3}$ being quasicircles, it is enough to show that $\xi$ can be extended as a quasiregular map in each of these domains. We extend it to $D_{1}$ and the other two are done in the same way. Let $\phi\colon\overline{D_{1}}\to\overline{\mathbb{D}}$ be the Böttcher coordinate of $\overline{D_{1}}$. Counting topological degrees, for example, we find that
$\phi\circ f^{2}=P_{4}\circ\phi\quad\mathrm{on}\quad\overline{D_{1}},$
where $P_{d}(z)=z^{d}$ as above. Therefore, $\xi\colon AB\to\partial D_{1}$ is conjugated to
$P_{4}(z)=z^{4}\colon\phi(AB)\to\mathbb{S}^{1}.$
Applying Lemma 8.1 to the conjugate of $\xi$ by $\phi$, we obtain a quasiregular extension into the unit disk. Conjugating it back by $\phi$, which is a conformal map, we obtain a quasiregular extension of $\xi$ into $D_{1}$, and hence a quasiregular extension of $\xi$ to $\widehat{\mathbb{C}}$, still denoted by $\xi$.
Finally, $\xi|_{\mathcal{J}(f)}$ cannot be extended to a global rational map. Indeed, if it extended to a rational map of $\widehat{\mathbb{C}}$, then, by uniqueness for analytic maps, it would have to equal to $f^{2}$ in $\Delta ABC$ and to the identity in $\widehat{\mathbb{C}}\backslash\left((\Delta ABC)\cup\sqcup_{i=1}^{3}D_{i}\right)$. This is a contradiction since $f^{2}$ is not equal to the identity.